byeeee:

What is the equation for a geometric sequence with a first term of 5 and a second term of −10? an = 5(−2)n − 1 an = 5(2)n − 1 an = 5(−15)n − 1 an = 5(15)n − 1

byeeee:

@dude

dude:

Geometric sequences are written as \(a_n=a_1\cdot r^{n-1}\) \(a_1\) is the first term \(r\) is the rate Do you have an idea on this?

byeeee:

no not really besides that 5 is a1

dude:

For this one we can work backwards, we know that the second term is -10 and the first term is 5, so... \(n\) is 2 \(a_n=-10\) and \(a_1=5\) \(a_n=a_1\cdot r^{n-1}\) \(-10=5\cdot r^{2-1}\) Combine the exponent \(-10=5\cdot r^{1}\) \(r^1\) can be re-written as just \(r\) \(-10=5\cdot r\) Divide by 5 on both sides \(-2=r\)

dude:

Now we have all values \(a_1=5\) \(r=-2\) So.. \(a_n=5(-2)^{n-1}\)

byeeee:

oh

byeeee:

ohh i see

byeeee:

Thank you